A 1,300 kg has an initial speed of 15 kph. How much force is required to accelerate the car to a speed of 120 kph in 10.0 s?
1
Expert's answer
2020-12-06T17:25:28-0500

"F=m\times a=m\times \frac{V-V_{0}}{t}=1300\times \frac{120-15}{10}=13650N"